3) Which one of the following groups cyclic? a) Z, ZA b) Z, Z c) Z, Z d) Z, Z6 e) Z.Z f) None

Please explain why each answer is wrong or correct

Thank you
Answer #1

3 a) Zzx 24 0,3 - 62 (0, 1), (0, 2) (1,0), (1, 1), (1, 2), (1,3 order (10,0)) = 0 order (0, 1) = 4 01) +(0,1)+(0,1) H0,1 = (0b - 12 Z₂ X ZG f(0, 1), 60,2), (0,3), (0,4), (0,5) (0, 0, 1, 1),(1,2),(13),(4) (1,2), (40) order (10,0)) = 1 ordes ((enl)= orDate (113) CWD (11) + (0,2) = 0 (1,99 = (1,1)+(1.3) = (2,0) (1, 1) = (1,1)+(2,0) (0, 1) (wwe= (1,) + (0,0) = (1/2) (LU = (1,24 X 76 grd (4,6)=2 #1 . 24 X 26 is not cyclic 6 .